OpenStudy (anonymous):

Find the similarity ratio of a prism with a surface area of 36 cm2 to a similar prism with a surface area of 400 cm2.

OpenStudy (nurali):

When you have a similarity ratio of some number x, then the larger has sides x times the small, surface areas x^2 times the larger, and volumes x^3 times the larger. 36/400 = 9/100 Since x^2 = 9/100, x = 3/10